# Impact of Exchange-Correlation Functionals on Predictions of Phonon Hydrodynamics: A Study of Fluorides, Chlorides, and Hydrides

### v2

@@ 4 Discussion
To reconcile the observation that the calculations of electronic band gap deviate significantly from experimental values, while better agreement is obtained for phonon properties, we examine a selection of modern XC functionals. For instance, Becke and Johnson (BJ) exchange potential was proposed to precisely regenerate the exact atomic exchange optimized effective potential (OEP) potential in terms of shape, resulting in accurately calculated electronic band gaps [tran2007band]. To assess band-gap accuracy using this functional, we evaluated LiF, NaCl, and LiH, obtaining 11.26, 7.81, and 4.80 eV, respectively, in agreement with measured values [roessler1967electronic, piacentini1976thermoreflectance, miyata1968optical, himpsel1978angle, kondo1988effect, plekhanov1998wannier]. However, this functional is less accurate in terms of structural calculations in comparison to PBESOL. The calculated lattice constants are 4.08 Å for LiF, 5.69 Å for NaCl, and 3.97 Å for LiH. A more accurate choice for structural calculations is the SCAN functional [sun2015scan]. The calculated lattice constants using the SCAN functional are 3.97 Å for LiF, 5.59 Å for NaCl, and 4.00 Å for LiH. Nevertheless, the calculated band gap using SCAN is less accurate than the BJ functional (LiF = 10.27 eV, NaCl = 6.10 eV, LiH = 3.58 eV). The HSE (Heyd-Scuseria-Ernzerhof) hybrid functional [heyd2003hybrid] is another functional that has been shown to improve electronic properties such as band gaps. We tested PBE with HSE in LiF and found that the band gap increased to 11.17 eV with an error of $\sim 18\%$. Since the HSE functional is not presently implemented within the ph.x package of Quantum ESPRESSO [giannozzi2009quantum], Phonopy [togo2023implementation, togo2023first] and Phono3py [togo2023implementation, togo2023first] are employed to calculate the second and third interatomic force constants, respectively. To ensure a fair comparison among exchange–correlation functionals, we assess the predicted thermal-transport regimes (see Fig. 7). Both HSE and SCAN trends fall between PBE and PBESOL.
@@
One possible reason phonon calculations remain weakly, but non-negligibly impacted, by the choice of functional and thereby ensure the reasonable accuracy of thermal calculations in contrast to electronic calculations, is the distinct impact of absolute total energies on electronic calculations and the influence of relative energy differences on phonon calculations [carbogno2022numerical]. Absolute total energies in electronic calculations suffer from inaccuracies due to the approximations made within the XC functionals. On the other hand, phonon calculations (i.e., force constants) depend on energy differences, which can be computed more reliably since any consistent error or systematic offset in energy calculations may cancel out, resulting in an accurate description of thermal properties. In other words, phonon properties depend on the shape of the interatomic potential energy surface and not its absolute value. Even so, the influence of exchange-correlation functionals on phonon properties is evident as demonstrated in the variability of the phonon hydrodynamics windows for all the materials studied in this work.
@@
It is insightful to consider the steps required to produce reliable first-principles calculations in the absence of experimental data. As a starting point, one can begin with materials with similar atomic compositions for which functionals have been developed and validated. In the absence of a well-established pseudopotential, one can obtain a preliminary pseudopotential using a pseudopotential generation code, such as those based on the projector-augmented wave (PAW) method. This preliminary pseudopotential can be refined through iterative calculations, comparing the results with available experimental and theoretical data in related systems, even if indirect. Convergence of DFT input parameters, such as k-point sampling and cutoff energy, is required to remove additional sources of uncertainty. Additionally, beyond the selection of pseudopotential and exchange-correlation functional, there is ongoing work to benchmark the large number of available DFT codes. For instance, Bosoni et al. [bosoni2024verify] proposed a protocol to enhance the reliability and precision of DFT methods and codes, providing a substantial reference dataset to foster improvements and ensure reproducibility in computational studies. The protocol leverages AiiDA workflows to streamline and ensure reproducibility in verifying DFT computations using a dataset based on all-electron methods spanning the entire periodic table [bosoni2024verify]. This effort can be interpreted as a way to measure numerical error (i.e., choice of algorithms and not the choice of XC functional) in DFT codes, which can be combined with error due to the functional approximations via error propagation approaches [parks2020uncertainty].
@@ 5 Conclusions
Using DFT, we calculated the electrical, mechanical, and thermal properties of NaF, LiF, KF, KCl, NaCl, LiH, NaH and KH using PBE, PBESOL, and LDA functionals. Phonon scattering rates were used to predict phonon hydrodynamics windows based on Guyer’s condition and solutions to the linearized BTE. The impact of isotopes on lattice thermal properties and phonon hydrodynamics windows was characterized. We find that GGA functionals, when compared to the LDA functionals, do not necessarily lead to more accurate predictions of the thermal conductivity. Overall, our calculations demonstrate good agreement with, when available, previous experimental and previous theoretical results, but non-negligible differences in the prediction of the phonon hydrodynamics window due to the choice of functional were observed.
CRediT authorship contribution statement
Jamal Abou Haibeh: Conceptualization (Equal); Data curation (Lead); Formal analysis (Lead); Investigation (Equal); Methodology (Equal); Validation (Equal); Visualization (Lead); Writing – original draft (Lead); Writing – review & editing (Equal). Samuel Huberman: Conceptualization (Equal); Data curation (Supporting); Formal analysis (Supporting); Funding acquisition (Lead); Investigation (Equal); Methodology (Equal); Project administration (Lead); Resources (Lead); Software (Lead); Supervision (Lead); Validation (Equal); Visualization (Supporting); Writing – original draft (Supporting); Writing – review & editing (Equal).
Declaration of competing interest
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
@@ Acknowledgements
Supplementary data
Supplementary data to this article can be found online at the supplementary material.
Data availability
The data that supports the findings of this study are available within the article and its supplementary material.

### v3

@@ 4 Discussion
To reconcile the observation that the calculations of electronic band gap deviate significantly from experimental values, while better agreement is obtained for phonon properties, we examine a selection of modern XC functionals. For instance, Becke and Johnson (BJ) exchange potential was proposed to precisely regenerate the exact atomic exchange optimized effective potential (OEP) in terms of shape, resulting in accurately calculated electronic band gaps [135]. To assess band-gap accuracy using this functional, we evaluated LiF, NaCl, and LiH, obtaining 11.26, 7.81, and 4.80 eV, respectively, in agreement with measured values [107, 98, 81, 52, 67, 100]. However, this functional is less accurate in terms of structural calculations in comparison to PBEsol. The calculated lattice constants are 4.08 Å for LiF, 5.69 Å for NaCl, and 3.97 Å for LiH. A more accurate choice for structural calculations is the SCAN functional [122]. The calculated lattice constants using the SCAN functional are 3.97 Å for LiF, 5.59 Å for NaCl, and 4.00 Å for LiH. Nevertheless, the calculated band gap using SCAN is less accurate than the BJ functional (LiF = 10.27 eV, NaCl = 6.10 eV, LiH = 3.58 eV). The HSE (Heyd-Scuseria-Ernzerhof) hybrid functional [50] is another functional that has been shown to improve electronic properties such as band gaps. We tested PBE with HSE in LiF and found that the band gap increased to 11.17 eV with an error of $\sim 18\%$. Since the HSE functional is not presently implemented within the ph.x package of Quantum ESPRESSO [40], Phonopy [129, 131] and Phono3py [129, 131] were employed to calculate the second and third interatomic force constants, respectively. To ensure a fair comparison among exchange–correlation functionals, we assess the predicted thermal-transport regimes (see figure 7). Both HSE and SCAN trends fall between PBE and PBEsol.
@@
One possible reason phonon calculations remain weakly, but non-negligibly impacted, by the choice of functional and thereby ensure the reasonable accuracy of thermal calculations in contrast to electronic calculations, is the distinct impact of absolute total energies on electronic calculations and the influence of relative energy differences on phonon calculations [21]. Absolute total energies in electronic calculations suffer from inaccuracies due to the approximations made within the XC functionals. On the other hand, phonon calculations (i.e., force constants) depend on energy differences, which can be computed more reliably since any consistent error or systematic offset in energy calculations may cancel out, resulting in an accurate description of thermal properties. In other words, phonon properties depend on the shape of the interatomic potential energy surface and not its absolute value. Even so, the influence of exchange-correlation functionals on phonon properties is evident as demonstrated in the variability of the phonon hydrodynamics windows for all the materials studied in this work.
@@
It is insightful to consider the steps required to produce reliable first-principles calculations in the absence of experimental data. As a starting point, one can begin with materials with similar atomic compositions for which functionals have been developed and validated. In the absence of a well-established pseudopotential, one can obtain a preliminary pseudopotential using a pseudopotential generation code, such as those based on the projector-augmented wave (PAW) method. This preliminary pseudopotential can be refined through iterative calculations, comparing the results with available experimental and theoretical data in related systems, even if indirect. Convergence of DFT input parameters, such as k-point sampling and cutoff energy, is required to remove additional sources of uncertainty. Additionally, beyond the selection of pseudopotential and exchange-correlation functional, there is ongoing work to benchmark the large number of available DFT codes. For instance, Bosoni et al [12] proposed a protocol to enhance the reliability and precision of DFT methods and codes, providing a substantial reference dataset to foster improvements and ensure reproducibility in computational studies. The protocol leverages AiiDA workflows to streamline and ensure reproducibility in verifying DFT computations using a dataset based on all-electron methods spanning the entire periodic table [12]. This effort can be interpreted as a way to measure numerical error (i.e., choice of algorithms and not the choice of XC functional) in DFT codes, which can be combined with error due to the functional approximations via error propagation approaches [91].
@@ 5 Conclusions
Using DFT, we calculated the electrical, mechanical, and thermal properties of NaF, LiF, KF, KCl, NaCl, LiH, NaH, and KH using PBE, PBEsol, and LDA functionals. Phonon scattering rates were used to predict phonon hydrodynamics windows based on Guyer’s condition and solutions to the linearized BTE. The impact of isotopes on lattice thermal properties and phonon hydrodynamics windows was characterized. We find that GGA functionals, when compared to the LDA functionals, do not necessarily lead to more accurate predictions of the thermal conductivity. Overall, our calculations demonstrate good agreement with, when available, previous experimental and previous theoretical results. We find non-negligible differences in the prediction of the phonon hydrodynamics window due to the choice of functional. Building on recent first-principles studies that explicitly include four-phonon scattering in BTE calculations [44, 127, 126], incorporating four-phonon processes in future work would further improve quantitative accuracy by refining the resulting average phonon scattering rates.
Data availability statement
All data that support the findings of this study are included within the article (and any supplementary files).
@@ Acknowledgements
Conflict of interest
The authors declare no financial competing interest.
Author contributions
Jamal Abou Haibeh   0009-0003-7167-4731
Conceptualization (Equal); Data curation (Lead); Formal analysis (Lead); Investigation (Equal); Methodology (Equal); Validation (Equal); Visualization (Lead); Writing – original draft (Lead); Writing – review & editing (Equal).
Samuel Huberman   0000-0003-0865-8096
Conceptualization (Equal); Data curation (Supporting); Formal analysis (Supporting); Funding acquisition (Lead); Investigation (Equal); Methodology (Equal); Project administration (Lead); Resources (Lead); Software (Lead); Supervision (Lead); Validation (Equal); Visualization (Supporting); Writing – original draft (Supporting); Writing – review & editing (Equal).